This exercise refers to a standard deck of playing cards. Assume that 5 cards are randomly chosen from the deck.
How many hands contain exactly 3 queens?

Answer by Edwin McCravy(20054)

(4 queens CHOOSE 3)×(48 non-queens CHOOSE 2) = (4C3)×(48C2) = 4×1128 = 4512

Edwin